### v1

# VLAP: Efficient Video-Language Alignment via Frame Prompting and Distilling for Video Question Answering

Abstract
In this work, we propose an efficient Video-Language Alignment via Frame-Prompting and Distilling (VLAP) network. Our VLAP model addresses both efficient frame sampling and effective cross-modal alignment in a unified way. In our VLAP network, we design a new learnable question-aware Frame-Prompter together with a new cross-modal distillation (QFormer-Distiller) module. Pre-trained large image-language models have shown promising results on problems such as visual question answering. However, how to efficiently and effectively sample image frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge. Compared with prior work, our VLAP model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up). Overall, our VLAP network outperforms (e.g. +4.6% on STAR Interaction and +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on VLEP with 4.2$\times$ speed up) the state-of-the-art methods on the video question-answering benchmarks.
1 Introduction
“If a picture is worth thousands of words, what is a video worth?” [32] Video watching is growing into a new social norm. Statistic shows YouTube has approximately $122$ million daily active users, based all over the world. Visitors spend on average $19$ minutes per day on YouTube. An average of close to 1 million hours of video are streamed by YouTube users each and every minute. As video data continue to grow through internet, video information retrieval becomes more and more demanding. Video data has tremendous capacity to store a vast variety of useful information. Compared to images, video understanding is difficult in the one extra temporal dimension. How to efficiently sample relevant frames from a video with the computing resource constraint remains a long-standing problem in video understanding research.
Figure 1: Our efficient Vision-Language Alignment via Frame-Promper and distilling (VLAP) model, at core, contains two new modules: a question-aware Frame-Prompter and a cross-modal QFormer-Distiller. It learns to extract the most question-related frames while keeping the inference latency low.
Recent advances in pre-trained large-scale language models [6, 42] have greatly boosted the performance of the vision-language models. Especially for image-language models [7, 1, 23, 46], many state-of-the-art works leverage pre-trained LLMs. These models  [9, 57] achieve excellent performance across visual-language tasks such as image captioning [9], visual question answering [57] and more. Inherently, many video-language models [62, 29] are built from these pre-trained image-language models. These image-based video-language models treat a video as a series of multi-channel images sampled randomly or uniformly [23]. While this strategy works well for short videos, for long videos or videos with non-uniform information distribution, random or uniform frame sampling may miss critical information. When it comes to video-language alignment, frame sampling efficiency and effectiveness goes hand-in-hand. One needs to not only reduce the number of sampled frames but also select frames that are most related to the input question. Previous work such as SeViLA[62] trains a separate keyframe localizer, which is not friendly for the real-time inference and introduce more parameters to the whole model. Besides video representation, cross-modality alignment while leveraging LLMs is another challenge. The critical problem lies in how to transfer video information to the LLM’s input domain.
Figure 2: Model Overview. Our VLAP model includes 4 sub-modules: the visual encoder, question-aware Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM. We encode the video frames through a frozen visual encoder. Then we train the Teacher-QFormer using all the frame features. After that, we train the Student-QFormer and Frame-Prompter end-to-end. Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a question-aware Frame-Prompter. Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer.
To address these challenges, we propose a new network, VLAP. Compared to the state-of-the-art video-language models [62, 9, 23], VLAP consists of a new question-aware Frame-Prompter together and a new QFormer-Distiller. To align image and language domains, we apply the cross-modal query-visual Transformer (QFormer) proposed in previous BLIP models [9, 23, 24] for cross-modal fusion. To improve the efficiency, we add a new distillation on top of the QFormer. We train our new Frame-Prompter and new QFormer-Distiller end-to-end. Our new question-aware Frame-Prompter learns to select the most relevant frames according to the input question text, and the QFormer-Distiller further reinforces the critical frames selection. To the best of our knowledge, we are the first to propose a Frame-Prompter and a distiller on top of the cross-modal alignment for video-language learning with pre-trained LLMs.
We validate our VLAP model on the Video Question Answering benchmark datasets. This includes the NExT-QA [52], STAR[50], How2QA [27], TVQA [21] and VLEP [22]. Our work outperforms previous strong SOTA methods across all the benchmark while reducing inference latency. Comparing with SOTA video-language model SeViLA [62], we reduce significant amount ($50\%$) of training parameters and inference latency (1.354.2×1.35-4.2$\times$1.35 - 4.2 × speed up), while improving the accuracy. We also performed ablation study on our question-aware Frame-Prompter and QFormer-Distiller. As shown in Table 4, both the question-aware Frame-Prompter (FP) and the QFormer-Distiller (QFD) play critical roles in making our method effective.
To summarize, our main contribution is a new question-aware Frame-Prompter and QFormer-Distiller component (trained from end-to-end) to efficiently and effectively select frames for video-language alignment learning.
2 Related Work
2.1 Visual-Language Alignment
Visual-Language Pre-training Vision-Language cross-modal pre-training has greatly improved over the past couple of years. Various network architectures and pre-training objectives have been proposed for different downstream tasks, including the dual-encoder architecture with image-text contrastive learning [38], the fusion-encoder architecture with image-text matching [25], and unified transformer architecture with masked language modeling [45]. These methods, along with others, focus on the ability to find image-text affinity [58], correlation [4], and/or completion [61], and need to pre-train the model end-to-end. To address the incompatibility with pre-trained unimodal models such as LLMs [6], recent works [23] proposed to train a QFormer to bridge the domain gap between two frozen pre-trained models. Inspired by its flexibility, more downstream tasks and applications have been proposed, including instruction-based image generation [51] and image question-answering [57].
While most of the previous work focus on image-text alignment, very few have discussed the extension to videos until most recently when temporal modeling starts to be included for better reasoning capabilities. HiTeA [59] jointly trains pairs in long and short view to capture temporal relations between moments and event. Smaug [28] introduces sparse image patch masks to reduce pre-training costs and improve cross-modal alignment. EgoVLPv2 [36] proposes cross-attention between the backbone encoders to improve both the pre-training efficiency and the downstream task performance. Yang et al. [56] models trajectory attention as a two-step attention operation to further emphasize word-related embeddings during the pre-training.
Our method leverages the pre-trained model during the video-language alignment, which naturally provides more capability to reasoning with less training cost.
Image-to-Video Transfer Learning
Due to the high cost of large video dataset collection, many works leverage successful pre-trained image models and transfer the knowledge to video task. Previous works such as  [3, 5, 16, 26] utilize a pretrained ViT [11] and aggregate the temporal image feature sequence using transformer block to adapt for video understanding task. For Video-Language task, many works turn to large-scale Vision-Language models as the starting point, such as CLIP [38], BLIP [24]. Many works choose to adapt a pre-trained CLIP model for text-to-video retrieval task, by either augmenting the frame-level spatial features with temporal information [53, 10], or manipulate the cross-modal similarity calculation to get better video-language alignment [31, 30, 12, 49]. Other works  [35, 37] focus on parameter efficient fine-tuning for video task by inserting trainable temporal modules into the pre-trained transformer architecture while keeping the rest of model frozen. Recent work SeViLA [62] proposes a language-aware frame localizer to sample relevant key frames from videos. It adopts a pre-trained Vision-Language model BLIP2 [23] as the frozen backbone and only the adapter QFormer is trained. In this paper, we propose a trainable question-aware Frame-Prompter and a QFormer-Distiller module. These techniques help to efficiently bridge the gap between image-language and video-language learning.
Video Question Answering
One major downstream task of Video-Language pre-training is Video Question Answering (VQA). Early works [2, 19, 60] often rely on human annotated datasets, while recent works [54, 64, 63] make better use of large-scale data from public. Along with the quick advances in VQA methods, A lot of benchmark datasets have also been introduced to standardize the model performance comparison, including NExT-QA [52], STAR [50], How2QA [27], TVQA [21], and VLEP [22]. We benchmark our network mainly on the Video Question and Answering task. It’s also the major task to test the effectiveness of the video-language alignment.
2.2 Knowledge Distillation
One of our key component is cross-modal distillation. Knowledge Distillation [17, 65, 47] is original proposed for small and efficient models to mimic the softened class distributions, features of large teachers. For multi-modalities, researchers explore how to utilize the prior knowledge between different modalities [15, 40, 41]. On video domain, knowledge distillation has been used for efficient video inference [33, 20], video captioning [34], video question answering [39]. Out of the supervised learning methods mentioned above, the idea of knowledge distillation has also been leveraged in many self-supervised methods for self-supervised video representation learning [44].
3 Method
Our VLAP model tackles the following challenges in large-scale Video-Language learning: how to sample question related frames and how to transfer visual information for pre-trained frozen LLMs. In this section, we describe our question-aware Frame-Prompter and QFormer-Distiller.
3.1 Model Architecture
As illustrated in Fig. 2, VLAP has four components: a pre-trained frozen large scale visual encoder $E_{v}$, a question-aware Frame-Prompter $F_{p}$, a QFormer-Distiller $Q_{d}$ (Querying Alignment Transformer [24, 23, 9] with distillation) to fuse and extract text-conditioned visual information, and a pre-trained frozen LLM. We train our Frame-Prompter, QFormer-Distiller and two other frozen modules end-to-end. Our training objective include a distillation loss $L_{\text{distill}}$ and a task loss $L_{\text{task}}$. For the multi-choice Video Question Answer task, the task loss is the cross-entropy loss.
3.2 Question-aware Frame-Prompter
Figure 3: Frame-Prompter. Here we show the details of our learnable Frame-Prompter. We design a learnable Frame-Prompter to sample the most text query-related frames. We encode the mean-pooled segment features. We then apply the Gumbel Softmax to compute the segment mask to guarantee the differentiability. The selected frames embedding then goes through the QFormer-Distiller. Here $B$ means batch size, $T$ means frame number, $N\times C$ means the frame feature sequences.
For the four-dimensional video data, it’s impractical and non-efficient to input all frames into a visual encoder model. Here we design a question-aware Frame-Prompter for efficient and effective frame sampling. It’s designed learn to attend to fewer but more question related frames. We start from the uniformly sampled video frames $\{f_{1},\dots,f_{T}\}$, $T$ is frame number. As shown in Fig. 3, these raw frames first go through the pre-trained visual encoder $E_{v}$,
$X=\{x_{i}|x_{i}=E_{v}(f_{i}),i\in[1,T]\},$ (1)
where $x_{i}$ is the visual feature extracted from raw frames and $T$ is the number of frames. We perform mean-pooling on the channels per-patch then encode the mean-pooled visual features onto an embedding using a fully-connected (FC) layer with a layer normalization (LN),
$\hat{x_{i}}=W_{2}*\operatorname{ReLU}(\mathrm{LN}(W_{1}*\operatorname{Mean}(x_{i}))),$ (2)
where $W_{1}$, $W_{2}$ are the learnable weights and $\operatorname{LN}$ is the layer normalization.
Our Frame-Prompter is a differentiable frame selecter. The differentiability is achieved through the Gumbel-Softmax [18]. We first convert a segment of concatenated frames feature $s_{k}=[x_{i},\dots,x_{S}]$ into a categorical distribution $\pi$ through the softmax operation,
$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(x_{j})},k,j\in[1,T/S]\right\},$ (3)
where $S$ is the number of frames of a segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
[…]
where $\mathcal{H}$ is the one-hot encoding operation. During training, we replace the argmax with the softmax for differentiability,
$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],$ (5)
where $\tau$ is a tunable temperature hyper-parameter.
3.3 Cross-Modal Distillation
In this section, we present our QFormer-Distiller, designed to further enforce the text conditioned visual information extraction. Video-language alignment performance, unlike image-language alignment, is closely related the selected frames. Meanwhile, to leverage powerful pre-trained LLMs, we need to transform the selected visual information to the LLM’s input domain. We adopt the QFormer proposed in the BLIP models [24, 23, 9] for the cross-modal transformation learning.
The original QFormer is designed for image-text fusion, and we add a distillation module to make it efficient for video-text fusion. Like the traditional distillation, our QFormer-Distiller includes a teacher-QFormer and a student-QFormer. We train the teacher-QFormer first (without the student-QFormer) on a wider receptive field. Then we train our student-QFormer to narrow down the most important frames. The student-teacher learning mechanism further encourages both the QFormer and the Frame-Prompter to learn to attend to the most relevant visual information given the input question text.
Specifically, during student learning process, both the teacher-QFormer and the student-QFormer will take in the video $X_{v}$ and question text $X_{q}$. The concatenated tokenized video and question text will go through the QFormer cross-attention layers:
$X^{\prime}=\operatorname{CrossAttn}(E_{t}(X_{q}),\operatorname{SelfAttn}(E_{v}(X_{v}),\mathbf{q})),$ (6)
where $E_{t}$ is a text tokenizer, $E_{v}$ is the visual tokenizer, $\mathbf{q}$ is the learnable query token. We utilize a decoder $D$ to transform the student-QFormer output, ensuring consistency and recoverability to the teacher’s feature. Then the distillation objective is defined by:
$L_{\text{distill}}=\operatorname{MSE}(D(X^{\prime}_{\text{student}}),X^{\prime}_{\text{teacher}}).$ (7)
We carefully design the decoder to be a simple Fully Connected (FC) layer with a layer normalization. We show through an ablation study (Sec. 5.3.2) that this combination works the best.
4 Experiment Setup
Here we introduce our implementation details, benchmark dataset and the baseline methods.
4.1 Implementation Details
For training, we conduct experiments with 8 × 40GB A100 GPUs. All the models in our experiments are trained using AdamW with cosine learning scheduler. For all the experiments, we use ViT-G(1B) [13] as the visual encoder and Flan-T5 XL (3B) [8] as large language model. For all the datasets, we use accuracy of choosing the right answer as the metric, and test on the validation dataset. For the temperature $\tau$ in the Frame-Prompter, it changes from 1 to 0.01 by $(0.01)^{(\text{current-step/total-training-step})}$ and is set as 0.01 during testing. And all inference (Infer.) time is evaluated on a single A100 GPU with batch size 4. Please refer supplementary for more training details.
4.2 Benchmark
To demonstrate the effectiveness of our proposed method on the video QA task, we compare our algorithm with the state-of-the-art (SOTA) methods on 5 datasets including 1 on video event prediction: 1) NExT-QA [52], a multi-choice VideoQA benchmark, with 3 types of questions: Causal (Why, How), Temporal (Previous/Next, Present), and Description (Binary, Location, Count and Other). 2) STAR[50], a multi-choice VideoQA benchmark for Situated Reasoning. It has four kinds of questions: Interaction, Sequence, Prediction, and Feasibility. 3) How2QA [27], a multi-choice VideoQA benchmark with 44k QA pairs for 22k 60-second clips selected from 9035 videos. 4) TVQA [21] is a large-scale video QA dataset with 152K questions along with 21k video clips from 460 hours of video. 5) VLEP [22] is a video event prediction benchmark, with 28,726 future event prediction cases. Following SeViLA [62], we formulate this task as a multi-choice Question Answering.
Method # Frames Temporal Causal Description Average Intro. Param. Infer. Time (ms/video)
Just Ask [54] (ICCV2021) 20 51.4 49.6 63.1 52.3 - -
All-in-One [43] (CVPR2023) 32 48.6 48.0 63.2 50.6 - -
MIST [14] (CVPR2023) 32 56.6 54.6 66.9 57.1 - -
HiTeA [59] (ICCV2023) 16 58.3 62.4 75.6 63.1 - -
InternVideo [48] (Dec 2022) 8 58.5 62.5 75.8 63.2 - -
VLAP (Ours) 1 66.5 69.3 78.0 70.5 188M 64
BLIP-2 [23] (ICML2023) 4 67.2 70.3 79.8 71.5 - -
VLAP (Ours) 2 70.2 71.6 79.4 72.8 188M 72
SeViLA [62] (NeurIPS2023) 4 67.7 72.1 82.2 73.4 376M 301
SeViLA [62] (NeurIPS2023) 8 67.0 73.8 81.8 73.8 376M 306
[…]
VLAP (Ours) 32 72.3 74.9 82.1 75.5 188M -
Table 1: Comparison Results on NExT-QA dataset. Here we measure the accuracy of choosing the right answer. Especially on Temporal and Causal type of questions, our VLAP (using only 4 frames) improves 3.3% and 1.7% respectively, compared with SeViLA. We use bold-face font to indicate the best results and underline on the second best using the same number of frames (brown box for 4 frames and blue box for 8 frames). VLAP using 2-frames only out-performs BLIP-2 using 4-frames by 1.3%. VLAP also achieves upto 3.04$\times$ speedup.
4.3 Baselines
We compare the performance of our VLAP model with several recent prominent models in the field. These include SeViLA [62], BLIP-2 [23], and InternVideo [48], within the context of fine-tuning scenarios. For fair comparison with SeViLA and BLIP-2, we employ the Vision Transformer-Global (ViT-G) as the visual encoder and Flan-T5-XL models as the Large Language Model (LLM). We follow SeViLA to adapt BLIP-2 for video-language training.
5 Results
Here we show both quantitative and qualitative (Sec. 5.2) comparison results between our VLAP and SOTA methods on Video QA and Video Event Prediction task (Sec. 5.1). Together we present an in-depth analysis on the results. Then we conduct ablation study (Sec. 5.3) on our proposed Frame-Prompter and QFormer-Distiller module and the design choices within each module.
5.1 Comparison Results on Video QA and Video Event Prediction
Overall, we demonstrate that the cross-modal key-frame selection matters. Our VLAP model out-performs strong SOTA methods across the 4 Video QA benchmark datasets and 1 Video Event Prediction while keeping the inference latency low. Especially, we highlight that our models’ performance stands out on temporal (including Causal, Interaction, Prediction) type of questions, NExT-QA Temporal (+3.3%, 3.04×3.04$\times$3.04 × speed up), NExT-QA Causal(+1.7%, 1.45×1.45$\times$1.45 × speed up), STAR Interaction (+4.6%, 3.04×3.04$\times$3.04 × speed up), STAR Prediction (+2.8%, 3.04×3.04$\times$3.04 × speed up), How2QA (+0.3%, 3.04×3.04$\times$3.04 × speed up), VLEP (+0.7% with 1.45×1.45$\times$1.45 × speed up, +0.3% with 4.18×4.18$\times$4.18 × speed up) and TVQA (+1.8%, 3.04×3.04$\times$3.04 × speed up).
On the NExT-QA [52] dataset, compared with the SOTA SeViLA [62] on 4-frame and 8-frame setting, our VLAP improves by 1.0% on average accuracy while achieving a 3.04×3.04$\times$3.04 × speed up (see Table 1).
Method (Frames Number) Interaction Sequence Prediction Feasibility Avg. Infer. Time (ms/video)
Flamingo-9B 4-shot  [1] (30) (NeurIPS2022) - - - - 42.8 -
All-in-One [43] (32) (CVPR2023) 47.5 50.8 47.7 44.0 47.5 -
MIST [14] (32) (CVPR2023) 55.5 54.2 54.2 44.4 51.1 -
InternVideo [48] (8) (Dec 2022) 62.7 65.6 54.9 51.9 58.7 -
BLIP-2 [23] (4) (ICML2023) 65.4 69.0 59.7 54.2 62.0 -
VLAP (2) (Ours) 65.0 65.4 62.2 58.8 62.9 72
SeViLA [62] (4) (NeurIPS2023) 63.7 70.4 63.1 62.4 64.9 301
[…]
Table 2: Comparison Results on STAR Video QA benchmark. For Interaction type of question, our VLAP improved 4.6%. On average, our VLAP out-performs the SOTA method by 2.2% when using 4 frames with 3.04$\times$ speed up. We bold the best results and underline the second best. We’d like to highlight that our VLAP using 2-frames out-performs BLIP-2 using 4-frames.
Method (# Frames) How2QA VLEP TVQA
FrozenBiLM (10)  [55] (NeurIPS2022) 81.5 - 57.5
InternVideo (8)  [48] (Dec 2022) 79.0 63.9 57.2
BLIP-2 (4)  [23] (ICML2023) 82.2 67.0 54.5
SeViLA (4)  [62] (NeurIPS2023) 83.6 68.9 61.6
VLAP (2) (Ours) 82.8 69.2 60.6
VLAP (4) (Ours) 83.9 69.6 63.4
Table 3: Comparison Results on How2QA, VLEP, and TVQA Video QA Benchmarks. VLAP improves performance over SeViLA by 1.8% with 3.04$\times$ speed up on TVQA dataset, 0.7% with 1.45×1.45$\times$1.45 × speed up on VLEP dataset, and 0.3% with 3.04×3.04$\times$3.04 × speed up on How2QA dataset at 4 frames setting. We’d like to highlight that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up.
We also test our VLAP model on the STAR [50] dataset. This dataset is challenging due to its diversified type of questions. As shown in Table 2, our VLAP model outperforms the several strong SOTA models on average by 2.2% with ${\bf 3.04}\times$ speed up. Especially on the STAR Interaction and Prediction type of questions, we outperform SeViLA [62] by 4.6% and 2.8%. This result further demonstrate the importance of key-frame selection. And our model’s advantages in extracting temporal and causal related key-frames.
We further test our model on the longer and larger scale video Video QA benchmark datasets: TVQA [21] and How2QA [27], as shown in Table 3. On the TVQA [21] dataset, our VLAP outperforms the SOTA method by 1.8% at the 4-frame setting. On How2QA, our VLAP improvement is 0.3% with 3.04×3.04$\times$3.04 × speed up compared with SeViLA [62]. This is partially due to the limited 32-frame teacher-QFormer training. On the other hand, compared with Blip-2 [23], VLAP outperforms by 1.7%. This difference again shows the key-frames selected by our VLAP model aligns better with the input question compared with uniform sampling. To test our algorithm’s capabilities, particularly for event prediction, we conduct an additional series of evaluations on VLEP [22] dataset. At the 4-frame setting, VLAP improves over SeViLA [62] by 0.7% with 1.45×1.45$\times$1.45 × speed up. it’s noteworthy that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up (Table 3).
Figure 4: Key-frame Selection Comparison Results (select 4 frames from 32 frames). We compare frames selected by our VLAP compared with that from the SOTA SeViLA [62] method. Across different type of questions, especially the Causal, Temporal type questions, keyframes selected by our network is more relevant and better related to the question.
5.2 Qualitative Results
We qualitatively compare the key-frames selected by our VLAP with the ones from SeViLA [62] on different type of questions. As shown in Figure 4, our network select more question-related key-frames across different question types (including Causal, Temporal and Description or Factual). In the first (col 1 row 1) example in Figure 4, our VLAP locates the frames that visibly show the “car and the dirt”, but the frames selected by SeViLA focuses on the “road”. In the second (col 2 row 1) example, we locate the frames with the “the man not on the bull at the end”, but the frames selected by SeViLA focuses on the “man on the bull at the end”. And in the fourth (col 2 row 2) example, we locate 3 frames with the “four people” according to the question, but non of the frames selected by SeViLA shows “four people”. For Temporal type of questions, our VLAP is also capable of selecting frames that closer to the action specified in the question. In the seventh (col 1 row 4) and eighth (col 2 row 4) example in Figure 4, we locate key-frames around the “the men on stage” and the “black dog” (vs. SeViLA has 2 frames focusing on the “brown dog”).
In addition, we qualitatively check the key-frames selected though our QFormer-Distiller. We show in Figure 5 training our QFormer-Distiller helps select the most question related frames. In the second example, we select frames that shows the “deep into the whole”. And in the third example, out of the 16 frames, we pick up the frames that have both the “man in grey and the stick falling”.
5.3 Ablation Study
Here we demonstrate the effectiveness of each proposed modules: the question-aware Frame-Prompter and QFormer-Distiller. We also validate the decoder design choice within the QFormer-Distiller.
Figure 5: QFormer-Distiller Results Visualization. Here we visualize the keyframes selected after cross-modal distillation. After distillation, we can select the most question-relevant frames even from 16 frames.
5.3.1 Frame-Prompter and QFormer-Distiller Ablation
Components STAR VLEP TVQA NExT-QA
base (BLIP-2) 62.0 67.0 54.5 71.5
base+QFD 64.9 68.6 62.2 73.5
base+QFD+FP 66.5 69.6 63.4 74.4
Table 4: Frame-Prompter and QFormer-Distiller Ablation Results. Across all four VideoQA datasets, we observe that both Question-aware Frame-Prompter (FP) and cross-modal QFormer-Distiller (QFD) contribute significantly to our final performance. We highlight that on STAR, adding our QFormer-Distiller improves the accuracy by 2.9%. Our Frame-Prompter (FP) further boost the accuracy by 1.6%.
Here we ablate our new QFormer-Distiller (QFD) and Frame-Prompter (FP). To summarize, our QFormer-Distiller and our Frame-Prompter each contributes 50% to the overall performance improvement across the 4 Video QA benchmarks. Specifically, our QFormer-Distiller boost performance by 2.9% and our Frame-Prompter by 1.6% on the STAR [50] dataset, as shown in Table 4. This shows both modules are critical to achieving the desirable performance.
5.3.2 QFormer-Distiller Decoder Ablation
We study the design choice of the decoder of the QFormer-Distiller. One of the key components is the decoder before computing the distillation loss. Design choices for this component includes the number of Fully Connected layer (FC) and where to put a layer normalization (LN). From Table 5, we show that the simple FC with a LN (after FC) works the best.
Frame Prompter T C D Average
FC 68.5 70.9 79.3 72.4
FC+LN 70.1 73.8 82.1 74.4
FC+LN+FC 69.7 73.1 81.6 74.1
Table 5: QFormer-Distiller Decoder Ablation on NExT-QA. We find that a simple Fully Connected layer (FC) with a Layer Normalization (LN) works best across Temporal (T), Causal (C), Description (D). It is efficient and effective.
The FC layer provides the necessary computational structure, while the LN layer stabilizes the learning process. This configuration balances the effectiveness and operational efficiency, making it a well-suited choice for our method.
6 Conclusion and Future Work
The advances in Large Language Models (LLMs) bring great boost to natural language understandings. New challenge emerges on how to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks. In this work, we present a new VLAP network with a new question-aware Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment. From our experiments, our VLAP out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions. We demonstrate that cross-modal keyframe selection is key to successful video-language alignment task execution. We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.

### v3

@@
-VLAP: Efficient Video-Language Alignment via Frame Prompting and Distilling for Video Question Answering
+ViLA: Efficient Video-Language Alignment for Video Question Answering
 Abstract
-In this work, we propose an efficient Video-Language Alignment via Frame-Prompting and Distilling (VLAP) network.
-Our VLAP model addresses both efficient frame sampling and effective cross-modal alignment in a unified way.
-In our VLAP network, we design a new learnable question-aware Frame-Prompter together with a new cross-modal distillation (QFormer-Distiller) module.
-Pre-trained large image-language models have shown promising results on problems such as visual question answering.
-However, how to efficiently and effectively sample image frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge.
-Compared with prior work, our VLAP model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up).
-Overall, our VLAP network outperforms (e.g.
-+4.6% on STAR Interaction and +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on VLEP with 4.2$\times$ speed up) the state-of-the-art methods on the video question-answering benchmarks.
+In this work, we propose an efficient Video-Language Alignment (ViLA) network.
+Our ViLA model addresses both efficient frame sampling and effective cross-modal alignment in a unified way.
+In our ViLA network, we design a new learnable text-guided Frame-Prompter together with a cross-modal distillation (QFormer-Distiller) module.
+Pre-trained large image-language models have shown promising results on problems such as visual question answering (VQA).
+However, how to efficiently and effectively sample video frames when adapting pre-trained large image-language model to video-language alignment is still the major challenge.
+Compared with prior work, our ViLA model demonstrates the capability of selecting key frames with critical contents, thus improving the video-language alignment accuracy while reducing the inference latency (+3.3% on NExT-QA Temporal with 3.0$\times$ speed up).
+Overall, our ViLA network outperforms the state-of-the-art methods on the video question-answering benchmarks: +4.6% on STAR Interaction, +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset with 4.2$\times$ speed-up.
 1 Introduction
-“If a picture is worth thousands of words, what is a video worth?” [32] Video watching is growing into a new social norm.
-Statistic shows YouTube has approximately $122$ million daily active users, based all over the world.
+“If a picture is worth thousands of words, what is a video worth?” [36] Video watching has become a new social norm.
+Statistics show YouTube has approximately $122$ million daily active users, based all over the world.
 Visitors spend on average $19$ minutes per day on YouTube.
 An average of close to 1 million hours of video are streamed by YouTube users each and every minute.
-As video data continue to grow through internet, video information retrieval becomes more and more demanding.
+As video data continue to grow through internet viewing, video information retrieval becomes more demanding.
 Video data has tremendous capacity to store a vast variety of useful information.
-Compared to images, video understanding is difficult in the one extra temporal dimension.
-How to efficiently sample relevant frames from a video with the computing resource constraint remains a long-standing problem in video understanding research.
-Figure 1: Our efficient Vision-Language Alignment via Frame-Promper and distilling (VLAP) model, at core, contains two new modules: a question-aware Frame-Prompter and a cross-modal QFormer-Distiller.
+Compared to image question answering (Q&A) problem, video QA is more challenging due to one extra temporal dimension.
+How to efficiently sample relevant frames from a video with the computing resource constraint remains a long-standing problem in video QA research.
+Figure 1: Our efficient Vision-Language Alignment (ViLA) model via Frame-Promper and distilling contains two new modules: a text-guided Frame-Prompter and a cross-modal QFormer-Distiller.
 It learns to extract the most question-related frames while keeping the inference latency low.
-Recent advances in pre-trained large-scale language models [6, 42] have greatly boosted the performance of the vision-language models.
-Especially for image-language models [7, 1, 23, 46], many state-of-the-art works leverage pre-trained LLMs.
-These models  [9, 57] achieve excellent performance across visual-language tasks such as image captioning [9], visual question answering [57] and more.
+Recent advances in pre-trained large-scale language models [6, 46] have greatly boosted the performance of the vision-language models, especially on cross-modality tasks.
+Many state-of-the-art image-language models [8, 1, 27, 49] leverage pre-trained LLMs.
+These models  [10, 60] achieve excellent performance on visual-language tasks such as image captioning [10], visual question answering [60] and more.
 Inherently, many video-language models [65, 33] are built from these pre-trained image-language models.
@@
 While this strategy works well for short videos, for long videos or videos with non-uniform information distribution, random or uniform frame sampling may miss critical information.
-When it comes to video-language alignment, frame sampling efficiency and effectiveness goes hand-in-hand.
+When it comes to video-language alignment, frame sampling efficiency and effectiveness go hand-in-hand.
 One needs to not only reduce the number of sampled frames but also select frames that are most related to the input question.
-Previous work such as SeViLA[62] trains a separate keyframe localizer, which is not friendly for the real-time inference and introduce more parameters to the whole model.
+Previous work such as SeViLA[65] trains a separate keyframe localizer, which is not friendly for the real-time inference and introduces more parameters to the whole model.
 Besides video representation, cross-modality alignment while leveraging LLMs is another challenge.
-The critical problem lies in how to transfer video information to the LLM’s input domain.
+The critical problem lies in how to efficiently transfer video information to the LLM’s input domain.
 Figure 2: Model Overview.
-Our VLAP model includes 4 sub-modules: the visual encoder, question-aware Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM.
+Our ViLA model includes 4 sub-modules: the visual encoder, text-supervised Frame-Prompter (FP), QFormer-Distiller (QFD), and a LLM.
 We encode the video frames through a frozen visual encoder.
@@
 After that, we train the Student-QFormer and Frame-Prompter end-to-end.
-Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a question-aware Frame-Prompter.
+Unlike the Teacher-QFormer, our Student-QFormer is trained with masked frames features from a text-supervised Frame-Prompter.
 Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer.
-To address these challenges, we propose a new network, VLAP.
-Compared to the state-of-the-art video-language models [62, 9, 23], VLAP consists of a new question-aware Frame-Prompter together and a new QFormer-Distiller.
-To align image and language domains, we apply the cross-modal query-visual Transformer (QFormer) proposed in previous BLIP models [9, 23, 24] for cross-modal fusion.
-To improve the efficiency, we add a new distillation on top of the QFormer.
-We train our new Frame-Prompter and new QFormer-Distiller end-to-end.
-Our new question-aware Frame-Prompter learns to select the most relevant frames according to the input question text, and the QFormer-Distiller further reinforces the critical frames selection.
-To the best of our knowledge, we are the first to propose a Frame-Prompter and a distiller on top of the cross-modal alignment for video-language learning with pre-trained LLMs.
-We validate our VLAP model on the Video Question Answering benchmark datasets.
+Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and fineuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
+Main Contributions: To address these challenges, we propose a new network, ViLA.
+Compared to the state-of-the-art video-language models [65, 10, 27], ViLA consists of a new Frame-Prompter together with a QFormer-Distiller.
+Our Frame-Prompter learns to select the most important frames influenced by the corresponding question text and and supervised by the VQA loss.
+Meanwhile, the Frame-Prompter is meticulously designed to keep lightweight so as to be efficient.
+To effectively and efficiently transfer video information to LLM input domain, we add a new distillation on top of the QFormer, named QFormer-Distiller.
+The QFormer is the cross-modal query-visual Transformer proposed in previous BLIP models [10, 27, 28] for cross-modal fusion.
+We train our Frame-Prompter and QFormer-Distiller end-to-end.
+The cross-modal temporal distiller teaches a smaller (i.e.
+fewer frames) QFormer.
+To the best of our knowledge, this work is the first to propose a Frame-Prompter and a distiller on top of the cross-modal alignment for video-language learning with pre-trained LLMs.
+We validate our ViLA model on the Video Question Answering benchmark datasets.
 This includes the NExT-QA [56], STAR[54], How2QA [31], TVQA [25] and VLEP [26].
-Our work outperforms previous strong SOTA methods across all the benchmark while reducing inference latency.
+Our work outperforms previous strong SOTA methods across all the benchmarks, while reducing inference latency.
 Comparing with SOTA video-language model SeViLA [65], we reduce significant amount ($50\%$) of training parameters and inference latency (1.354.2×1.35-4.2$\times$1.35 - 4.2 × speed up), while improving the accuracy.
-We also performed ablation study on our question-aware Frame-Prompter and QFormer-Distiller.
-As shown in Table 4, both the question-aware Frame-Prompter (FP) and the QFormer-Distiller (QFD) play critical roles in making our method effective.
-To summarize, our main contribution is a new question-aware Frame-Prompter and QFormer-Distiller component (trained from end-to-end) to efficiently and effectively select frames for video-language alignment learning.
+We also performed ablation study on our text-guided Frame-Prompter and QFormer-Distiller.
+As shown in Table 4, both the text-guided Frame-Prompter and the QFormer-Distiller play critical roles in making our method effective.
+To sum up, the key novelty is a new text-guided Frame-Prompter and question-relevant QFormer-Distiller (trained from end-to-end).
+The Frame-Prompter enhances the efficiency, while the later bolsters the effectiveness.
+Together they optimize the selection of frames for video-language alignment learning.
 2 Related Work
 2.1 Visual-Language Alignment
-Visual-Language Pre-training Vision-Language cross-modal pre-training has greatly improved over the past couple of years.
+Visual-Language Pre-training
+Vision-Language cross-modal pre-training has greatly improved over the past couple of years.
 Various network architectures and pre-training objectives have been proposed for different downstream tasks, including the dual-encoder architecture with image-text contrastive learning [42], the fusion-encoder architecture with image-text matching [29], and unified transformer architecture with masked language modeling [50].
@@
 EgoVLPv2 [40] proposes cross-attention between the backbone encoders to improve both the pre-training efficiency and the downstream task performance.
-Yang et al.
-[56] models trajectory attention as a two-step attention operation to further emphasize word-related embeddings during the pre-training.
-Our method leverages the pre-trained model during the video-language alignment, which naturally provides more capability to reasoning with less training cost.
+Our method can leverages the pre-trained model during the video-language alignment, which naturally provides reasoning capability with less training cost.
 Image-to-Video Transfer Learning
@@
 Recent work SeViLA [65] proposes a language-aware frame localizer to sample relevant key frames from videos.
-It adopts a pre-trained Vision-Language model BLIP2 [23] as the frozen backbone and only the adapter QFormer is trained.
-In this paper, we propose a trainable question-aware Frame-Prompter and a QFormer-Distiller module.
+In this paper, we propose a trainable text-guided Frame-Prompter and a QFormer-Distiller module, which help focus more on the important temporal and spatial information from the 2D frames.
 These techniques help to efficiently bridge the gap between image-language and video-language learning.
@@
 Out of the supervised learning methods mentioned above, the idea of knowledge distillation has also been leveraged in many self-supervised methods for self-supervised video representation learning [48].
+2.3 Frame Selection for Video QA
+Early approaches in Video QA relied heavily on dense sampling methods, where frames are extracted at a fixed interval throughout the video.
+While straightforward, such methods can lead to excessive computational costs and memory requirements without significantly improving performance.
+Zhang et al.
+[68] proposed a more selective strategy, using attention mechanisms to identify key frames that are more likely to contain information relevant to the question.
+Following this, adaptive frame sampling methods  [13, 7] have gained popularity.
+These methods aim to dynamically select frames based on the content’s relevance to the question, thus optimizing the trade-off between computational efficiency and answer accuracy [13].
 3 Method
-Our VLAP model tackles the following challenges in large-scale Video-Language learning: how to sample question related frames and how to transfer visual information for pre-trained frozen LLMs.
-In this section, we describe our question-aware Frame-Prompter and QFormer-Distiller.
+Our ViLA model tackles the following challenges in large-scale Video-Language learning: how to sample question related frames and how to efficiently transfer video information for pre-trained frozen LLMs.
 3.1 Model Architecture
 As illustrated in Fig.
-2, VLAP has four components: a pre-trained frozen large scale visual encoder $E_{v}$, a question-aware Frame-Prompter $F_{p}$, a QFormer-Distiller $Q_{d}$ (Querying Alignment Transformer [24, 23, 9] with distillation) to fuse and extract text-conditioned visual information, and a pre-trained frozen LLM.
+2, ViLA has four components: a pre-trained frozen large scale visual encoder $E_{v}$, a Frame-Prompter $F_{p}$, a QFormer-Distiller $Q_{d}$ (Querying Alignment Transformer [28, 27, 10] with distillation) to fuse and extract text-conditioned visual information, and a pre-trained frozen LLM.
 We train our Frame-Prompter, QFormer-Distiller and two other frozen modules end-to-end.
-Our training objective include a distillation loss $L_{\text{distill}}$ and a task loss $L_{\text{task}}$.
-For the multi-choice Video Question Answer task, the task loss is the cross-entropy loss.
-3.2 Question-aware Frame-Prompter
-Figure 3: Frame-Prompter.
-Here we show the details of our learnable Frame-Prompter.
-We design a learnable Frame-Prompter to sample the most text query-related frames.
-We encode the mean-pooled segment features.
+Our training objective includes a distillation loss $L_{\text{distill}}$ and a task loss $L_{\text{vqa}}$.
+For the multi-choice Video Question Answer task, the task loss is the cross-entropy.
+3.2 Text-guided Frame-Prompter Learning
+Figure 3: Text-guided Frame-Prompter.
+Here we show the details of our learnable text-guided Frame-Prompter.
+We design a learnable Frame-Prompter to sample the most text query-related frames, with two design choises (a and b).
+We choose design (a) for diversified temporal sampling.
+We first encode the mean-pooled segment features.
 We then apply the Gumbel Softmax to compute the segment mask to guarantee the differentiability.
@@
 Here $B$ means batch size, $T$ means frame number, $N\times C$ means the frame feature sequences.
+The Frame-Prompter is learned with the text-supervised gradient.
+When VQA loss is applied, the input question text-related gradient further flows to the Frame-Prompter.
+The question text-related gradient guides the Frame-Prompter to select the most critical frames.
 For the four-dimensional video data, it’s impractical and non-efficient to input all frames into a visual encoder model.
-Here we design a question-aware Frame-Prompter for efficient and effective frame sampling.
-It’s designed learn to attend to fewer but more question related frames.
+Here we design a text-guided Frame-Prompter for efficient and effective frame sampling.
+It is designed to learn attending to fewer but more question-related frames.
 We start from the uniformly sampled video frames $\{f_{1},\dots,f_{T}\}$, $T$ is frame number.
@@
 3, these raw frames first go through the pre-trained visual encoder $E_{v}$,
-$X=\{x_{i}|x_{i}=E_{v}(f_{i}),i\in[1,T]\},$ (1)
+$X=\{x_{i}|x_{i}=E_{v}(f_{i}),i\in[1,T]\},\vspace{-2mm}$ (1)
 where $x_{i}$ is the visual feature extracted from raw frames and $T$ is the number of frames.
-We perform mean-pooling on the channels per-patch then encode the mean-pooled visual features onto an embedding using a fully-connected (FC) layer with a layer normalization (LN),
-$\hat{x_{i}}=W_{2}*\operatorname{ReLU}(\mathrm{LN}(W_{1}*\operatorname{Mean}(x_{i}))),$ (2)
+We perform mean-pooling on the channels per-patch.
+The Frame-Prompter input shape is $(B,T,N,C)$, B batch size, T temporal, $(N,C)$ as the frame feature dimension.
+Mean-pool averages over the $C$ channel.
+After mean-pool, the feature dimension becomes $(B,T,N)$.
+Then we divide the feature into S segments $(B,S,T//S\cdot N)$, with $(T//S\cdot N)$ as the segment feature dimension.
+Then the segmented features goes through FC layer to project to $(B,S,T//S)$ dimension, ready for the gumbel-softmax computation.
+For free-form frame sampling Fig.
+3(b), after mean-pool, we use FC layer to transfer $(B,T,N)$ to $(B,T,2)$.
+Then we encode the mean-pooled visual features onto an embedding using a fully-connected (FC) layer with a layer normalization (LN),
+$\hat{x_{i}}=W_{2}*\operatorname{ReLU}(\mathrm{LN}(W_{1}*\operatorname{Mean}(x_{i}))),\vspace{-2mm}$ (2)
 where $W_{1}$, $W_{2}$ are the learnable weights and $\operatorname{LN}$ is the layer normalization.
 Our Frame-Prompter is a differentiable frame selecter.
+This allows the text-supervised gradients from the QFormer to guide the learning our frame selecter via backpropagation.
 The differentiability is achieved through the Gumbel-Softmax [21].
-We first convert a segment of concatenated frames feature $s_{k}=[x_{i},\dots,x_{S}]$ into a categorical distribution $\pi$ through the softmax operation,
-$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(x_{j})},k,j\in[1,T/S]\right\},$ (3)
+We have two choices (shown in Fig.
+3 a and b) for frame selection before the Frame-Prompter: a) strategic sampling (like: segmented uniform sampling) and b) free-form sampling.
+We choose option (a) to force diverse temporal sampling.
+We first convert a segment of concatenated frames feature $s_{k}=[\hat{x_{i}},\dots,\hat{x}_{T//S}]$ into a categorical distribution $\pi$ through the softmax operation,
+$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(s_{j})},k,j\in[1,T/S]\right\},\vspace{-2mm}$ (3)
 where $S$ is the number of frames of a segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
-[…]
+$M_{S}=\mathcal{H}\left(\arg\max_{k}\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
 where $\mathcal{H}$ is the one-hot encoding operation. During training, we replace the argmax with the softmax for differentiability,
-$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],$ (5)
-where $\tau$ is a tunable temperature hyper-parameter.
+$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],\vspace{-2mm}$ (5)
+where $\tau$ is a tunable temperature. Full mask is $M=\{M_{i},\cdots,M_{T//S}\}$. We apply mask $M$ to input frames and then conduct CrossAttention to obtain text guidance by
+$X_{LLM}=\operatorname{CrossAttn}(X\cdot M,X_{t}),$ (6)
+where $X_{t}$ is the text information.
+This CrossAttention can help the Frame-Prompter learn to choose rich text frames.
+Finally, we use the task loss below to supervise the learning and choose the frame selection to answer our question:
+$L_{\text{vqa}}=\operatorname{MSE}(LLM(X_{LLM},X_{t}),QA_{\text{answer}}).\vspace{-2mm}$ (7)
 3.3 Cross-Modal Distillation
-In this section, we present our QFormer-Distiller, designed to further enforce the text conditioned visual information extraction.
+In this section, we present in detail our cross-modal distillation module, designed to selectively transfer video information ready for LLM consumption.
 Video-language alignment performance, unlike image-language alignment, is closely related the selected frames.
@@
 We train the teacher-QFormer first (without the student-QFormer) on a wider receptive field.
-Then we train our student-QFormer to narrow down the most important frames.
 The student-teacher learning mechanism further encourages both the QFormer and the Frame-Prompter to learn to attend to the most relevant visual information given the input question text.
+We study how QFormer-Distiller affects Frame-Prompter (Sec.
+4.3.1).
+And we find our FP+QFD 4-frames(Acc.74.4%) model’s key-frame selection overlaps 92.3% with the key-frame selected by the FP 8-frames(Acc.74.1%) model.
 Specifically, during student learning process, both the teacher-QFormer and the student-QFormer will take in the video $X_{v}$ and question text $X_{q}$.
 The concatenated tokenized video and question text will go through the QFormer cross-attention layers:
-$X^{\prime}=\operatorname{CrossAttn}(E_{t}(X_{q}),\operatorname{SelfAttn}(E_{v}(X_{v}),\mathbf{q})),$ (6)
+$X^{\prime}=\operatorname{CrossAttn}(E_{t}(X_{q}),\operatorname{SelfAttn}(E_{v}(X_{v}),\mathbf{q})),\vspace{-2mm}$ (8)
 where $E_{t}$ is a text tokenizer, $E_{v}$ is the visual tokenizer, $\mathbf{q}$ is the learnable query token.
 We utilize a decoder $D$ to transform the student-QFormer output, ensuring consistency and recoverability to the teacher’s feature.
+This supervision bolsters the effectiveness and performance of the model with fewer frames.
 Then the distillation objective is defined by:
-$L_{\text{distill}}=\operatorname{MSE}(D(X^{\prime}_{\text{student}}),X^{\prime}_{\text{teacher}}).$ (7)
-We carefully design the decoder to be a simple Fully Connected (FC) layer with a layer normalization. We show through an ablation study (Sec. 5.3.2) that this combination works the best.
-4 Experiment Setup
-Here we introduce our implementation details, benchmark dataset and the baseline methods.
-4.1 Implementation Details
+$L_{\text{distill}}=\operatorname{MSE}(D(X^{\prime}_{\text{student}}),X^{\prime}_{\text{teacher}}).\vspace{-2mm}$ (9)
+We carefully design the decoder to be a simple Fully Connected (FC) layer with a layer normalization.
+We show through an ablation study (Sec.
+4.3.2) that this combination works the best.
+Meanwhile, our FP+QFD 4-frames boosts accuracy by 1.8% compared with FP 4-frames model and the key-frame selected overlaps 56.8% with that of the FP 4-frames model.
+This helps verify our QFD’s ability to both boost performance and enhance the key-frame selection.
+4 Experiments
+4.1 Implementation Settings
 For training, we conduct experiments with 8 × 40GB A100 GPUs.
@@
 Please refer supplementary for more training details.
-4.2 Benchmark
+4.1.1 Benchmark:
 To demonstrate the effectiveness of our proposed method on the video QA task, we compare our algorithm with the state-of-the-art (SOTA) methods on 5 datasets including 1 on video event prediction: 1) NExT-QA [56], a multi-choice VideoQA benchmark, with 3 types of questions: Causal (Why, How), Temporal (Previous/Next, Present), and Description (Binary, Location, Count and Other).
@@
 InternVideo [52] (Dec 2022) 8 58.5 62.5 75.8 63.2 - -
-VLAP (Ours) 1 66.5 69.3 78.0 70.5 188M 64
+ViLA (Ours) 1 66.5 69.3 78.0 70.5 188M 64
 BLIP-2 [27] (ICML2023) 4 67.2 70.3 79.8 71.5 - -
-VLAP (Ours) 2 70.2 71.6 79.4 72.8 188M 72
+ViLA (Ours) 2 70.2 71.6 79.4 72.8 188M 72
 SeViLA [65] (NeurIPS2023) 4 67.7 72.1 82.2 73.4 376M 301
@@
 […]
-VLAP (Ours) 32 72.3 74.9 82.1 75.5 188M -
+ViLA+LoRA (Ours) 4 71.4 74.5 80.3 75.1 188M 208
+ViLA (Ours) 32 71.8 75.3 82.1 75.6 188M -
 Table 1: Comparison Results on NExT-QA dataset.
 Here we measure the accuracy of choosing the right answer.
-Especially on Temporal and Causal type of questions, our VLAP (using only 4 frames) improves 3.3% and 1.7% respectively, compared with SeViLA.
+Especially on Temporal and Causal type of questions, our ViLA (using only 4 frames) improves 3.3% and 1.7% respectively, compared with SeViLA.
 We use bold-face font to indicate the best results and underline on the second best using the same number of frames (brown box for 4 frames and blue box for 8 frames).
-VLAP using 2-frames only out-performs BLIP-2 using 4-frames by 1.3%.
-VLAP also achieves upto 3.04$\times$ speedup.
-4.3 Baselines
-We compare the performance of our VLAP model with several recent prominent models in the field.
+ViLA using 2-frames only out-performs BLIP-2 using 4-frames by 1.3%.
+ViLA also achieves upto 3.04$\times$ speedup.
+It needs to be noted that our ViLA achieves 75.1% average accuracy with only 4 frames when we finetune LLM with LoRA [20].
+4.1.2 Baselines:
+We compare the performance of our ViLA model with several recent prominent models in the field.
 These include SeViLA [65], BLIP-2 [27], and InternVideo [52], within the context of fine-tuning scenarios.
-For fair comparison with SeViLA and BLIP-2, we employ the Vision Transformer-Global (ViT-G) as the visual encoder and Flan-T5-XL models as the Large Language Model (LLM).
-We follow SeViLA to adapt BLIP-2 for video-language training.
-5 Results
+For fair comparison with SeViLA and BLIP-2, we employ the Vision Transformer-Global (ViT-G) as the visual encoder and Flan-T5-XL models as the Large Language Model.
+4.2 Results
 Here we show both quantitative and qualitative (Sec.
-5.2) comparison results between our VLAP and SOTA methods on Video QA and Video Event Prediction task (Sec.
-5.1).
+4.2.2) comparison results between our ViLA and SOTA methods on Video QA and Video Event Prediction task (Sec.
+4.2.1).
 Together we present an in-depth analysis on the results.
 Then we conduct ablation study (Sec.
-5.3) on our proposed Frame-Prompter and QFormer-Distiller module and the design choices within each module.
-5.1 Comparison Results on Video QA and Video Event Prediction
+4.3) on our proposed Frame-Prompter and QFormer-Distiller module and the design choices within each module.
+4.2.1 Comparison Results on Video QA and Video Event Prediction:
 Overall, we demonstrate that the cross-modal key-frame selection matters.
-Our VLAP model out-performs strong SOTA methods across the 4 Video QA benchmark datasets and 1 Video Event Prediction while keeping the inference latency low.
+Our ViLA model out-performs strong SOTA methods across the 4 Video QA benchmark datasets and 1 Video Event Prediction while keeping the inference latency low.
 Especially, we highlight that our models’ performance stands out on temporal (including Causal, Interaction, Prediction) type of questions, NExT-QA Temporal (+3.3%, 3.04×3.04$\times$3.04 × speed up), NExT-QA Causal(+1.7%, 1.45×1.45$\times$1.45 × speed up), STAR Interaction (+4.6%, 3.04×3.04$\times$3.04 × speed up), STAR Prediction (+2.8%, 3.04×3.04$\times$3.04 × speed up), How2QA (+0.3%, 3.04×3.04$\times$3.04 × speed up), VLEP (+0.7% with 1.45×1.45$\times$1.45 × speed up, +0.3% with 4.18×4.18$\times$4.18 × speed up) and TVQA (+1.8%, 3.04×3.04$\times$3.04 × speed up).
-On the NExT-QA [52] dataset, compared with the SOTA SeViLA [62] on 4-frame and 8-frame setting, our VLAP improves by 1.0% on average accuracy while achieving a 3.04×3.04$\times$3.04 × speed up (see Table 1).
+On the NExT-QA [56] dataset, compared with the SOTA SeViLA [65] on 4-frame and 8-frame setting, our ViLA improves by 1.0% on average accuracy while achieving a ${\bf 3.04}\times$ speed up (see Table 1).
 Method (Frames Number) Interaction Sequence Prediction Feasibility Avg. Infer. Time (ms/video)
@@
 BLIP-2 [27] (4) (ICML2023) 65.4 69.0 59.7 54.2 62.0 -
-VLAP (2) (Ours) 65.0 65.4 62.2 58.8 62.9 72
+ViLA (2) (Ours) 65.0 65.4 62.2 58.8 62.9 72
 SeViLA [65] (4) (NeurIPS2023) 63.7 70.4 63.1 62.4 64.9 301
@@
 Table 2: Comparison Results on STAR Video QA benchmark.
-For Interaction type of question, our VLAP improved 4.6%.
-On average, our VLAP out-performs the SOTA method by 2.2% when using 4 frames with 3.04$\times$ speed up.
-We bold the best results and underline the second best.
-We’d like to highlight that our VLAP using 2-frames out-performs BLIP-2 using 4-frames.
-Method (# Frames) How2QA VLEP TVQA
-FrozenBiLM (10)  [55] (NeurIPS2022) 81.5 - 57.5
-InternVideo (8)  [48] (Dec 2022) 79.0 63.9 57.2
-BLIP-2 (4)  [23] (ICML2023) 82.2 67.0 54.5
-SeViLA (4)  [62] (NeurIPS2023) 83.6 68.9 61.6
-VLAP (2) (Ours) 82.8 69.2 60.6
-VLAP (4) (Ours) 83.9 69.6 63.4
+For Interaction type of question, our ViLA improved 4.6%.
+On average, our ViLA out-performs the SOTA method by 2.2% when using 4 frames with 3.04$\times$ speed up.
+Note that our ViLA using 2-frames out-performs BLIP-2 using 4-frames.
+Method Frames Numbers How2QA VLEP TVQA
+FrozenBiLM  [59] (NeurIPS2022) 10 81.5 - 57.5
+InternVideo  [52] (Dec 2022) 8 79.0 63.9 57.2
+BLIP-2  [27] (ICML2023) 4 82.2 67.0 54.5
+SeViLA  [65] (NeurIPS2023) 4 83.6 68.9 61.6
+ViLA (Ours) 2 82.8 69.2 60.6
+ViLA (Ours) 4 83.9 69.6 63.4
 Table 3: Comparison Results on How2QA, VLEP, and TVQA Video QA Benchmarks.
-VLAP improves performance over SeViLA by 1.8% with 3.04$\times$ speed up on TVQA dataset, 0.7% with 1.45×1.45$\times$1.45 × speed up on VLEP dataset, and 0.3% with 3.04×3.04$\times$3.04 × speed up on How2QA dataset at 4 frames setting.
-We’d like to highlight that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up.
-We also test our VLAP model on the STAR [50] dataset.
+ViLA improves performance over SeViLA by 1.8% with 3.04$\times$ speed up on TVQA dataset, 0.7% with 1.45×1.45$\times$1.45 × speed up on VLEP dataset, and 0.3% with 3.04×3.04$\times$3.04 × speed up on How2QA dataset at 4 frames setting.
+Ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up.
+We also test our ViLA model on the STAR [54] dataset.
 This dataset is challenging due to its diversified type of questions.
-As shown in Table 2, our VLAP model outperforms the several strong SOTA models on average by 2.2% with ${\bf 3.04}\times$ speed up.
+As shown in Table 2, our ViLA model outperforms the several strong SOTA models on average by 2.2% with ${\bf 3.04}\times$ speed up.
 Especially on the STAR Interaction and Prediction type of questions, we outperform SeViLA [65] by 4.6% and 2.8%.
@@
 And our model’s advantages in extracting temporal and causal related key-frames.
-We further test our model on the longer and larger scale video Video QA benchmark datasets: TVQA [21] and How2QA [27], as shown in Table 3.
-On the TVQA [21] dataset, our VLAP outperforms the SOTA method by 1.8% at the 4-frame setting.
-On How2QA, our VLAP improvement is 0.3% with 3.04×3.04$\times$3.04 × speed up compared with SeViLA [62].
+We further test our model on the longer and larger-scale Video QA benchmark datasets: TVQA [25] and How2QA [31], as shown in Table 3.
+On the TVQA [25] dataset, our ViLA outperforms the SOTA method by 1.8% at the 4-frame setting.
+On How2QA, our ViLA improvement is 0.3% with 3.04×3.04$\times$3.04 × speed up compared with SeViLA [65].
 This is partially due to the limited 32-frame teacher-QFormer training.
-On the other hand, compared with Blip-2 [23], VLAP outperforms by 1.7%.
-This difference again shows the key-frames selected by our VLAP model aligns better with the input question compared with uniform sampling.
+On the other hand, compared with Blip-2 [27], ViLA outperforms by 1.7%.
+This difference again shows the key-frames selected by our ViLA model aligns better with the input question compared with uniform sampling.
 To test our algorithm’s capabilities, particularly for event prediction, we conduct an additional series of evaluations on VLEP [26] dataset.
-At the 4-frame setting, VLAP improves over SeViLA [62] by 0.7% with 1.45×1.45$\times$1.45 × speed up.
-it’s noteworthy that ours 2-frames out-perform SeViLA 4-frames on VLEP by 0.3% with ${\bf 4.2}\times$ speed up (Table 3).
+At the 4-frame setting, ViLA improves over SeViLA [65] by 0.7% with 1.45×1.45$\times$1.45 × speed up.
+it’s noteworthy that ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset by 0.3% with ${\bf 4.2}\times$ speed up (Table 3).
 Figure 4: Key-frame Selection Comparison Results (select 4 frames from 32 frames).
-We compare frames selected by our VLAP compared with that from the SOTA SeViLA [62] method.
+We compare frames selected by our ViLA compared with that from the SOTA SeViLA [65] method.
 Across different type of questions, especially the Causal, Temporal type questions, keyframes selected by our network is more relevant and better related to the question.
-5.2 Qualitative Results
-We qualitatively compare the key-frames selected by our VLAP with the ones from SeViLA [62] on different type of questions.
+4.2.2 Qualitative Results:
+We qualitatively compare the key-frames selected by our ViLA with the ones from SeViLA [65] on different type of questions.
 As shown in Figure 4, our network select more question-related key-frames across different question types (including Causal, Temporal and Description or Factual).
-In the first (col 1 row 1) example in Figure 4, our VLAP locates the frames that visibly show the “car and the dirt”, but the frames selected by SeViLA focuses on the “road”.
+In the first (col 1 row 1) example in Figure 4, our ViLA locates the frames that visibly show the “car and the dirt”, but the frames selected by SeViLA focuses on the “road”.
 In the second (col 2 row 1) example, we locate the frames with the “the man not on the bull at the end”, but the frames selected by SeViLA focuses on the “man on the bull at the end”.
-And in the fourth (col 2 row 2) example, we locate 3 frames with the “four people” according to the question, but non of the frames selected by SeViLA shows “four people”.
-For Temporal type of questions, our VLAP is also capable of selecting frames that closer to the action specified in the question.
+And in the fourth (col 2 row 2) example, we locate 3 frames with the “four people” according to the question, but none of the frames selected by SeViLA shows “four people”.
+For Temporal type of questions, our ViLA is also capable of selecting frames that are closer to the action specified in the question.
 In the seventh (col 1 row 4) and eighth (col 2 row 4) example in Figure 4, we locate key-frames around the “the men on stage” and the “black dog” (vs.
@@
 4.3 Ablation Study
-Here we demonstrate the effectiveness of each proposed modules: the question-aware Frame-Prompter and QFormer-Distiller. We also validate the decoder design choice within the QFormer-Distiller.
+Here we demonstrate the effectiveness of each proposed module: the text-guided Frame-Prompter and the question-relevant QFormer-Distiller.
+We also validate the decoder design choice within the QFormer-Distiller.
 Figure 5: QFormer-Distiller Results Visualization.
@@
 After distillation, we can select the most question-relevant frames even from 16 frames.
-5.3.1 Frame-Prompter and QFormer-Distiller Ablation
+4.3.1 Frame-Prompter and QFormer-Distiller Ablation:
 Components STAR VLEP TVQA NExT-QA
 base (BLIP-2) 62.0 67.0 54.5 71.5
-base+QFD 64.9 68.6 62.2 73.5
-base+QFD+FP 66.5 69.6 63.4 74.4
+base+QFormer-Distiller 64.9 68.6 62.2 73.5
+base+Frame-Prompter 65.3 68.8 62.7 73.6
+base+QFormer-Distiller+Frame-Prompter 66.5 69.6 63.4 74.4
 Table 4: Frame-Prompter and QFormer-Distiller Ablation Results.
-Across all four VideoQA datasets, we observe that both Question-aware Frame-Prompter (FP) and cross-modal QFormer-Distiller (QFD) contribute significantly to our final performance.
+Across all four VideoQA datasets, we observe that both Text-aware Frame-Prompter and cross-modal QFormer-Distiller contribute significantly to our final performance.
 We highlight that on STAR, adding our QFormer-Distiller improves the accuracy by 2.9%.
-Our Frame-Prompter (FP) further boost the accuracy by 1.6%.
+Our Frame-Prompter further boost the accuracy by 1.6%.
 Here we ablate our new QFormer-Distiller (QFD) and Frame-Prompter (FP).
@@
 This shows both modules are critical to achieving the desirable performance.
-5.3.2 QFormer-Distiller Decoder Ablation
+We further explore how QFormer-Distiller (QFD) affect Frame-Prompter (FP).
+We find our FP+QFD 4-frames(Acc.74.4%) model’s key-frame selection overlaps 92.3% with the key-frame selected by the FP 8-frames(Acc.74.1%) model.
+Meanwhile, our FP+QFD 4-frames boosts accuracy by 1.8% compared with FP 4-frames model and the key-frame selected overlaps 56.8% with that of the FP 4-frames model.
+This helps verify our QFD’s ability to both boost performance and enhance the key-frame selection.
+4.3.2 QFormer-Distiller Decoder Ablation:
 We study the design choice of the decoder of the QFormer-Distiller.
@@
 From Table 5, we show that the simple FC with a LN (after FC) works the best.
-Frame Prompter T C D Average
+Frame Prompter Decoder Temporal Causal Description Average
 FC 68.5 70.9 79.3 72.4
 FC+LN 70.1 73.8 82.1 74.4
-FC+LN+FC 69.7 73.1 81.6 74.1
+FC+LN+GELU+FC 69.7 73.1 81.6 74.1
 Table 5: QFormer-Distiller Decoder Ablation on NExT-QA.
-We find that a simple Fully Connected layer (FC) with a Layer Normalization (LN) works best across Temporal (T), Causal (C), Description (D).
+We find that a simple Fully Connected layer (FC) with a Layer Normalization (LN) works best across Temporal, Causal, Description.
 It is efficient and effective.
+GELU is activation function.
 The FC layer provides the necessary computational structure, while the LN layer stabilizes the learning process.
 This configuration balances the effectiveness and operational efficiency, making it a well-suited choice for our method.
-6 Conclusion and Future Work
-The advances in Large Language Models (LLMs) bring great boost to natural language understandings.
-New challenge emerges on how to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks.
-In this work, we present a new VLAP network with a new question-aware Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment.
-From our experiments, our VLAP out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions.
-We demonstrate that cross-modal keyframe selection is key to successful video-language alignment task execution.
+4.4 More Discussion
+More recent works [24, 65] have shown powerful networks such as LLMs learns most of the data distribution during the pre-training stage.
+This is one of the major reason why prompt-learning is very effective.
+Our work is inspired from the prompt-learning.
+We hope to leverage LLM through proper visual prompting without affecting the generalization ability of the LLM, and this won’t affect LLMs’ original ability on language tasks.
+However, for the VQA task itself, optimal performance in this task often necessitates training the Language Model (LLM).
+Therefore, we conduct an ablation study on NExT-QA that fine-tunes the LLM by using LoRA during the training.
+Our ViLA achieves 75.1% average accuracy with only 4 frames.
+This demonstrates that our ViLA has a strong adaptation ability.
+Furthermore, to evaluate the generalization of our ViLA, we replace Flan-T5 with llama on NExT-QA, baseline with llama (68.6%) vs.
+ViLA with llama ( 72.7%), which shows that ViLA can adapt to different LLMs.
+5 Conclusion, Limitation and Future Work
+How to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks?
+In this work, we present a new ViLA network with a new text-guided Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment.
+From our experiments, our ViLA out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions.
+For the limitation, we don’t have a change to evaluate on LLMs larger than 13b due to the resource constraints.
 We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.

### v4

@@
 Abstract
-In this work, we propose an efficient Video-Language Alignment (ViLA) network.
+We propose an efficient Video-Language Alignment (ViLA) network.
 Our ViLA model addresses both efficient frame sampling and effective cross-modal alignment in a unified way.
@@
 Overall, our ViLA network outperforms the state-of-the-art methods on the video question-answering benchmarks: +4.6% on STAR Interaction, +2.2% on STAR average with 3.0$\times$ speed up, ours 2-frames out-perform SeViLA 4-frames on the VLEP dataset with 4.2$\times$ speed-up.
+Code will be available at https://github.com/xijun-cs/ViLA.
 1 Introduction
@@
 Finally, the input question text and QFormer transformed visual features go through a frozen large language model to generate the answer.
-Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and fineuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
+Our network supports both leveraging LLM through proper visual prompting without affecting the original LLM (Frozen) ability on language tasks and finetuning LLMs(LoRA) simultaneously to get optimal performance on specific tasks.
 Main Contributions: To address these challenges, we propose a new network, ViLA.
@@
 We benchmark our network mainly on the Video Question and Answering task.
-It’s also the major task to test the effectiveness of the video-language alignment.
 2.2 Knowledge Distillation
@@
 We first convert a segment of concatenated frames feature $s_{k}=[\hat{x_{i}},\dots,\hat{x}_{T//S}]$ into a categorical distribution $\pi$ through the softmax operation,
-$\pi=\left\{p_{k}\mid p_{k}=\frac{\exp(s_{k})}{\sum_{j=1}^{T/S}\exp(s_{j})},k,j\in[1,T/S]\right\},\vspace{-2mm}$ (3)
-where $S$ is the number of frames of a segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
-$M_{S}=\mathcal{H}\left(\arg\max_{k}\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
+$\pi=\left\{p_{k}\mid\frac{\exp(s_{i})}{\sum_{j=1}^{S}\exp(s_{j})},k\in[1,T/S],i,j\in[1,S]\right\},\vspace{-2mm}$ (3)
+where $S$ is the frame number of one segment. Then we compute our segment mask $M_{S}$ using the segment probability $p_{k}$ and the $g_{k}$ is sampled from the $\text{Gumbel}(0,1)$ distribution,
+$M_{k}=\mathcal{H}\left(\arg\max\left[g_{k}+\log p_{k}\right]\right)k\in[1,T/S],\vspace{-2mm}$ (4)
 where $\mathcal{H}$ is the one-hot encoding operation. During training, we replace the argmax with the softmax for differentiability,
-$\frac{\exp((\log p_{k}+g_{k})/\tau)}{\sum_{j=1}^{T/S}\exp((\log p_{j}+g_{j})/\tau)},k,j\in[1,T/S],\vspace{-2mm}$ (5)
+$\frac{\exp((\log p_{i}+g_{i})/\tau)}{\sum_{j=1}^{S}\exp((\log p_{j}+g_{j})/\tau)},i,j\in[1,S],\vspace{-2mm}$ (5)
 where $\tau$ is a tunable temperature. Full mask is $M=\{M_{i},\cdots,M_{T//S}\}$. We apply mask $M$ to input frames and then conduct CrossAttention to obtain text guidance by
@@
 5 Conclusion, Limitation and Future Work
-How to properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks?
+How do we properly ingest visual information to LLMs to utilize its capability effectively in cross-modal tasks?
 In this work, we present a new ViLA network with a new text-guided Frame-Prompter to smartly sample important frames, together with a cross-modal temporal distillation (QFormer-Distiller) for efficient and effective video-language alignment.
-From our experiments, our ViLA out-performs SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal (including Temporal, Causal, Interaction and etc.) type of questions.
-For the limitation, we don’t have a change to evaluate on LLMs larger than 13b due to the resource constraints.
-We plan to continue research on the design of our Frame-Prompter, especially on long video-language alignment.
+From our experiments, our ViLA outperforms SOTA methods on four video question answering benchmarks and one video event prediction benchmark, especially on the temporal and interaction type of questions.
+We demonstrate that cross-modal keyframe selection is key to successful video-language alignment task execution.
+Due to resource constraints, we only evaluate on LLMs with the number of parameters not larger than 13 billions.
+We plan to continue research on the design of our Frame-Prompter, especially on video-language alignment for long video segments.